At the beginning of the year, a stock's share price was $60. In March, it had a 3:5 reverse split and in June it paid $5 dividends. It closed the year with the share price of $45. What was the return for the stock for the year?

-46.67%

33.33%

-16.67%

38.89%
Answer #1

Return for the year:

= (Ending value+Dividends-Beginning value)/Beginning value

= ($45×5+$5×5-$60×3)/($60×3)

= ($225+$25-$180)/$180

= $70/$180

= 38.89%

Hence, correct option is 38.89%
Answer #2

To calculate the return for the stock for the year, we need to consider the change in stock price and any dividends paid.

Given information:

  • Beginning share price: $60

  • Reverse split in March: 3:5

  • Dividends paid in June: $5

  • Year-end share price: $45

To calculate the return for the stock for the year, we'll follow these steps:

  1. Adjust for the reverse split: After the reverse split, the new share price can be calculated as (old share price / reverse split ratio). In this case, the reverse split ratio is 3:5, which means the new share price is (60 / (3/5)) = $100.

  2. Calculate the total return without dividends: Total return without dividends = (Year-end share price - Beginning share price) / Beginning share price * 100 Total return without dividends = (45 - 60) / 60 * 100 Total return without dividends = -15 / 60 * 100 Total return without dividends = -25%

  3. Calculate the total return with dividends: Total return with dividends = (Year-end share price + Dividends paid - Beginning share price) / Beginning share price * 100 Total return with dividends = (45 + 5 - 60) / 60 * 100 Total return with dividends = -10 / 60 * 100 Total return with dividends = -16.67%

Therefore, the return for the stock for the year, considering both the change in stock price and dividends, is approximately -16.67%. Thus, the correct answer is -16.67%.


answered by: mervetokaz